# Bridging Fairness and Explainability: Can Input-Based Explanations Promote Fairness in Hate Speech Detection?

### v1

@@ Appendix F Fairness Correlations in Explanation-Debiased Models
Appendix G Faithfulness as an Indicator of Bias Detection Ability
What factors influence the reliability of explanations in detecting bias? In this section, we examine the relationship between explanation faithfulness and their ability to identify bias, reflected by fairness correlation scores in RQ1. We assess the faithfulness of explanation methods using two perturbation-based metrics: comprehensiveness and sufficiency AOPC (Area Over the Perturbation Curve; DeYoung et al., 2020), computed by masking 5%, 10%, 20%, and 50% of the input tokens. For substitution, we use the [MASK] token in BERT and the [PAD] token in Qwen3. Higher comprehensiveness and lower sufficiency scores indicate more faithful explanations.
Table 6: Faithfulness results of different explanation methods on BERT and Qwen3 models.
BERT Qwen3
Attention 4.50 3.20 61.70 10.34 17.20 14.17
Attn rollout 4.37 3.11 58.60 9.04 15.70 1.86
Attn flow 4.01 3.46 46.73 10.57 16.82 4.38
Grad L2 4.82 2.99 49.57 12.30 16.09 31.80
Grad mean 0.77 6.16 6.37 11.41 17.50 6.40
DeepLift L2 4.72 3.09 50.18 12.44 16.17 31.08
DeepLift mean 1.68 5.75 6.04 10.78 18.69 6.11
IxG L2 4.89 2.95 48.99 12.35 16.27 31.53
IxG mean 7.44 1.70 29.69 9.99 18.82 17.07
IntGrad L2 4.81 3.02 56.60 12.33 16.86 24.52
IntGrad mean 10.68 -0.36 45.31 14.21 16.12 4.46
Occlusion 13.16 -0.90 62.14 20.05 13.73 26.98
Occlusion abs 0.79 0.56 65.77 22.48 20.36 26.41
KernelSHAP 5.99 2.30 20.99 11.49 17.86 1.86
Appendix H LLM Usage

### v2

@@ Appendix I Faithfulness as an Indicator of Bias Detection Ability
Appendix M Limitations and Future Work
Our study has several limitations that we acknowledge and aim to address in future work.
First, as the first quantitative investigation of this topic, our study focuses solely on hate speech detection and uses a limited set of experimental setups. Although the results are consistent across these setups and preliminary experiments (Appendix H) suggest good generalization across tasks, models and sensitive token vocabulary, broader validation is still needed. Future work could extend this evaluation to additional domains and high-stakes applications.
Second, several findings are derived under the specific experimental setups used in this work. For instance, in RQ2, we conclude that the proposed attribution-based metrics are not reliable fairness indicators. However, it remains possible that other metrics could be effective. Similarly, our fairness-balanced metric in RQ3 may not be the optimal validation strategy in all settings. As it is infeasible to exhaustively enumerate and evaluate all potential configurations, we believe our conclusions nonetheless offer valuable guidance and highlight important methodological considerations for the community.
Third, our work focuses on evaluating standalone explanation-based strategies for improving fairness. Ensembles of multiple explanation methods, or hybrid approaches that combine explanation techniques with established debiasing methods, may yield better outcomes. Additionally, incorporating human oversight may further enhance the effectiveness and robustness of explanation-based fairness auditing. Our preliminary experiments show promising results in using hybrid debiasing techniques K, and demonstrates the possibility for human fairness auditing based on explanations J. Based on that, we believe that a systematic investigation of such hybrid or human-in-the-loop approaches represents an interesting avenue for future work.
Fourth, we do not identify any explanation method that consistently outperforms others across all research questions, which prevents us from offering a single recommendation. We therefore encourage future researchers to choose explanation methods that align with their specific tasks and constraints. Future work could further investigate why certain methods are better suited to particular settings and, ideally, develop practical guidelines for selecting effective methods without requiring extensive empirical comparisons.
Finally, although we consider both group and individual fairness, this work provides a more in-depth analysis of individual fairness (in RQ1 and RQ2), driven by the conceptual alignment between input-based explanations and individual fairness notions. We encourage future work to more thoroughly examine how explanation methods relate to group fairness.
Appendix N LLM Usage